Ajax Manufacturing - Budget

Ajax Manufacturing produces a single product, which takes 8.0 pounds of direct materials per unit produced. Assume that it is currently at the end of the first quarter of the year, and there are 50,000 pounds of material on hand. The company’s policy is to maintain an end-of-quarter inventory of materials equal to 25 percent of the following quarter’s material requirements for production. How many units of product were produced in the first quarter of the year? Under the assumption that production will increase by 10 percent in the second quarter, what are the direct materials requirements (in pounds) for planned production in the second quarter?
Answer #1

Part 1:

Closing balance DM   inventory - Q1 50000pounds25%
Total DM inventory for Q150000 x(100/25) =                      200,000pounds100%
DM inventory used per unit  8 pounds
Number of units produced in Q1200,000 / 8 = 25,000units


Part 2:


Expected units  produced in Q225,00 x 110% =                  27,500units
DM required for production27,500 x 8 pounds =  220,000pounds
Add: 25% closing balance requirement220,000 x 25% =  55,000pounds
Total DM requirements for   Q2 220,000 + 55,000 =  275,000pounds


answered by: HJN